OpenStudy (anonymous):

Simplify (radical) 7m^5 over (radical) 11

OpenStudy (anonymous):

you got it fruit haha

OpenStudy (anonymous):

\[\sqrt{7x^5}/\sqrt{11}\]\[(\sqrt{7x^5}\sqrt{11})/11\]

OpenStudy (anonymous):

i don't get it...

OpenStudy (anonymous):

what's the answer

OpenStudy (anonymous):

\[\sqrt{77x^5}/11\]

OpenStudy (anonymous):

oh, because you multiply the 11 and 7?

OpenStudy (anonymous):

thank you